OpenStudy (anonymous):

Part 1: Change 3x + 2y = 10 to slope-intercept form. You must show all work. (2 points) Part 2: State the slope and y-intercept (2 points)

OpenStudy (anonymous):

I did the work and narrowed it down to 2 answers

OpenStudy (anonymous):

y = (-3/2) x + 5

OpenStudy (anonymous):

y = (3/2) x -5

ganeshie8 (ganeshie8):

first one is correct

OpenStudy (anonymous):

3x + 2y = 10 2y= -3x+ 10 y= -(3/2)x +5

OpenStudy (anonymous):

Thank you